Let A and B be disjoint events and assume that Pr(A) = Pr(B) and Pr(AUB) = 0.8. Then Pr(A) =

Select one:

a. 0.4

b. 0.1

c. 0.2

d. 0.3


1
Expert's answer
2021-06-14T14:08:06-0400

If two events "A"  and "B" are disjoint, then


"Pr(A \cup B) = Pr(A) + Pr(B)."

Hence


"Pr(A)=Pr(A \cup B)- Pr(B)"

Given "Pr(A)=Pr(B)"

"Pr(A)=Pr(A \cup B)- Pr(A)"

"Pr(A)=0.5Pr(A \cup B)"

"Pr(A)=0.5(0.8)=0.4"

a. 0.4